Are the following pairs of lines parallel, perpendicular or neither. Please show me your work.
x-2y=3
2x+y=3
My answer: x -2y = 3
x + x -2y = -x + 3
-2y/-2 = -x/-2 + 3/-2
y = 1/2x -3/2
and 2x + y = 3
-2x + 2x + y = -2x + 3
y = -2x + 3
My guess is neither?
Help, please.

Answer by richwmiller(17219)

ax+by=c
slope=-a/b
x-2y=3
slope -1/-2=1/2
2x+y=3
slope -2/1=-2
1/2*-2=-1 therefore
perpendicular